OpenStudy (anonymous):

4. A membership to Movie Night Movie Club costs $10 + $2 per movie. Write the linear equation for this problem in y=mx+b form. 5. A membership to Movie Night Movie Club costs $10 + $2 per movie. Make a T-table for the equation for this problem using the x-values 0, 5, 10. Type in the ordered pairs you get as your answer 6. A membership 2 Movie Night Movie Club costs $10 + $2 per movie. Use the graph you made for the equation for this problem 2 approximate the answer 2 the following question u joined the Movie Night Movie Club and rented eight movies. How much did this cost according 2 ur graph

OpenStudy (anonymous):

4) y=2x+10

OpenStudy (anonymous):

x|y 1,20 2,40 3,60 4,80 5,100 and so on...

OpenStudy (anonymous):

hope that helps :)

OpenStudy (anonymous):

What's number 6?

OpenStudy (anonymous):

these are the answers to 4) and 5a)

OpenStudy (anonymous):

Do you know 6 though

OpenStudy (anonymous):

yes. it would be y=2(8)+10 y=16+10 y=26

OpenStudy (anonymous):

yes. it would be y=2(8)+10 y=16+10 y=26

OpenStudy (anonymous):

Thanks & can you help me with the rest of my questions that I've posted? Please